# DynaNoise: Dynamic Probabilistic Noise Injection for Defending Against Membership Inference Attacks

### v1

@@ Abstract.
Membership Inference Attacks (MIAs) pose a significant risk to the privacy of training datasets by exploiting subtle differences in model outputs to determine whether a particular data sample was used during training. These attacks can compromise sensitive information, especially in domains such as healthcare and finance, where data privacy is paramount. Traditional mitigation techniques, such as static differential privacy, rely on injecting a fixed amount of noise during training or inference. However, this approach often leads to a detrimental trade-off: the noise may be insufficient to counter sophisticated attacks or, when increased, may substantially degrade model performance. In this paper, we present DynaNoise, an adaptive approach that dynamically modulates noise injection based on query sensitivity. Our approach performs sensitivity analysis using measures such as Shannon entropy to evaluate the risk associated with each query and adjusts the noise variance accordingly. A probabilistic smoothing step is then applied to re-normalize the perturbed outputs, ensuring that the model maintains high accuracy while effectively obfuscating membership signals. We further propose an empirical metric, the Membership Inference Defense Privacy–Utility Trade-off (MIDPUT), which quantifies the balance between reducing attack success rates and preserving the target model’s accuracy. Our extensive evaluation on several benchmark datasets demonstrates that DynaNoise not only significantly reduces MIA success rates but also achieves up to a four-fold improvement in the MIDPUT metric compared to the state-of-the-art. Moreover, DynaNoise maintains competitive model accuracy while imposing only marginal inference overhead, highlighting its potential as an effective and efficient privacy defense against MIAs.
@@ 1. Introduction
Machine learning has revolutionized many domains by leveraging vast amounts of data to achieve impressive performance (Islam et al., 2024; Forough et al., 2023; Rayed et al., 2024; Zhang et al., 2019; Forough et al., 2022; Sahu et al., 2023). However, this success comes at a cost, where sensitive information from training datasets may be inadvertently memorized, posing serious privacy risks (Schwarzschild et al., 2024; El Mestari et al., 2024; Yang et al., 2024; Ye, 2024). For example, in a scenario where a hospital deploys a predictive model to diagnose diseases; if an attacker can determine whether a patient’s record was part of the training set, it may reveal that the patient has visited the hospital, thereby compromising their privacy. Similarly, in finance, membership leakage could expose clients’ investment histories. Such risks underscore the need for robust defenses against privacy attacks.
Membership Inference Attack (MIA) (Shokri et al., 2017), as shown in Figure 1, exploits subtle differences in a model’s output behavior to determine whether a specific data record was used during training, thereby threatening the confidentiality of individual data points (Shokri et al., 2017). Conventional defenses, such as differential privacy, introduce a fixed level of noise during training or inference to obscure membership information (Abadi et al., 2016). While these approaches offer formal privacy guarantees, they force an inherent trade-off between privacy and utility. In many practical settings, uniformly adding noise can significantly degrade model performance, yet reducing the noise level may leave the model susceptible to advanced membership inference attacks.
Figure 1. An illustration of how MIA works.
To address these challenges, we introduce DynaNoise, an adaptive noise injection approach that modulates the privacy noise dynamically based on query sensitivity. Our method first assesses the risk of each query through sensitivity analysis, by utilizing metrics such as Shannon entropy (Mageed and Zhang, 2022), and then adjusts the noise variance accordingly. A subsequent probabilistic smoothing step is applied to re-normalize the perturbed outputs, ensuring that the model retains high predictive accuracy while effectively obfuscating membership signals. Another key innovation of our approach is the introduction of a novel empirical metric, the Membership Inference Defense Privacy–Utility Trade-off (MIDPUT), which quantitatively captures the balance between the reduction in attack success rates and the preservation of model accuracy. Our experimental results on CIFAR-10, ImageNet-10, and SST-2 datasets demonstrate that DynaNoise not only significantly reduces membership inference attack success rates but also achieves up to a four-fold improvement in the MIDPUT metric compared to existing defenses such as SELENA.
@@
    Proposal of a novel defense mechanism against membership inference attacks called DynaNoise, that dynamically adjusts the noise level based on query sensitivity, thereby providing stronger privacy protection with minimal impact on target model accuracy.
@@
    Proposal of a new empirical metric called Membership Inference Defense Privacy–Utility Trade-off (MIDPUT), that quantifies the trade-off between the reduction in attack success rates and the loss in model performance. This metric enables a more detailed and precise evaluation of privacy defenses.
@@
    Comprehensive evaluation through extensive experiments on multiple benchmark datasets and models using different system parameters.
@@ 2.1. Membership Inference Attacks (MIAs)
Shadow training-based attacks. Shokri et al. (Shokri et al., 2017) introduced one of the earliest frameworks for membership inference attacks. In their approach, the adversary constructs multiple shadow models that mimic the behavior of the target model by training them on data drawn from a distribution similar to that of the target. The outputs of these shadow models on inputs with known membership statuses are then used to train an attack classifier that distinguishes between members and non-members. Although this method can achieve high attack accuracy, its effectiveness relies on access to a dataset closely resembling the target model’s training data.
Salem et al. (Salem et al., 2018) build upon this framework by demonstrating that a single shadow model can often suffice to perform effective membership inference. By reducing the number of shadow models, their method decreases the overall computational cost and query burden, making the attack more practical. However, the approach still assumes that the adversary can obtain or generate data from a similar distribution as the target model, which may not always be realistic.
Long et al. (Long et al., 2020) further refine shadow training-based attacks by optimizing both the construction of the shadow models and the design of the attack classifier. Their improvements reduce the number of queries needed and enhance the robustness of the attack, even when the adversary’s resources are limited. This work underscores that shadow training-based methods remain potent, though they continue to depend on the availability of representative auxiliary data.
Metric-based attacks. In contrast to shadow training-based methods, metric-based attacks infer membership directly from the target model’s output by evaluating specific statistical metrics. Yeom et al. (Yeom et al., 2018) propose a loss-based attack that exploits the tendency of models to incur lower loss on training samples compared to non-training samples. By setting an appropriate loss threshold, the adversary can effectively distinguish between members and non-members. While this approach is straightforward and does not require training additional models, its effectiveness is highly sensitive to the chosen threshold and may vary across different models.
Similarly, Song et al. (Song and Mittal, 2021) explore confidence-based attacks, which leverage the observation that models often yield higher prediction confidence for training samples. In this method, membership is inferred by comparing the maximum confidence score against a threshold, which can be either fixed or adjusted on a per-class basis. Although this technique is simple and requires little auxiliary information, it is also sensitive to the threshold setting, thereby affecting its overall robustness.
@@ 2.2. Membership Inference Defenses
Abadi et al. (Abadi et al., 2016) propose DP-SGD, a defense that integrates differential privacy directly into the training process by clipping per-example gradients and adding calibrated Gaussian noise. Their method employs a moments accountant to tightly track the cumulative privacy loss, thereby offering formal, provable privacy guarantees. This approach ensures that the entire model adheres to a specified privacy budget. However, because DP-SGD uses a static noise injection scheme that does not adapt to the sensitivity of individual queries, the large noise required to meet a strict privacy budget often leads to a significant drop in accuracy. Consequently, the overall utility of the trained model may suffer, especially in complex or high-dimensional tasks.
In another work, Nasr et al. (Nasr et al., 2018) introduce adversarial regularization, an approach that augments the training process with a min-max game. In their approach, an auxiliary attack model is trained concurrently with the main classifier to predict membership information, while the classifier is regularized to minimize the success of this inference attack. This empirical strategy has been shown to reduce the effectiveness of membership inference attacks with only a modest loss in utility. Nevertheless, because it does not offer formal differential privacy guarantees and its performance is sensitive to the tuning of adversarial hyperparameters, the method may not be robust across different datasets and deployment scenarios.
@@
Finally, Tang et al. (Tang et al., 2022) take an ensemble-based approach with SELENA, which trains multiple sub-models on overlapping subsets of the training data and then distills their outputs into a single prediction through a self-distillation process. This adaptive inference strategy selectively aggregates predictions from sub-models that have not seen the queried sample, thereby enhancing the privacy-utility trade-off. Despite its advantages, the ensemble inference process introduces moderate computational overhead due to the requirement of running multiple sub-models concurrently in training phase, which can be a drawback in resource-constrained settings.
@@
Defense Method Year Approach/Mechanism Privacy Guarantee Utility Impact Comp. Overhead
@@
MemGuard (Jia et al., 2019) 2019 Post-processing: adversarial noise added to confidence score vectors Empirical Minimal accuracy drop Low
@@
DynaNoise (This work) 2025 Adaptive noise injection based on query sensitivity (sensitivity analysis, dynamic noise variance modulation, probabilistic smoothing) Empirical Minimal accuracy drop Low
@@ 3. Problem Statement
Membership inference attacks exploit subtle discrepancies between a model’s outputs on training data and those on unseen data, thereby threatening the privacy of individuals whose information is used during training. Traditional privacy-preserving techniques, such as static differential privacy, address this risk by adding a fixed amount of noise to the model outputs. However, the uniform application of noise fails to account for the heterogeneous sensitivity of different queries. For low-risk queries, the excessive noise degrades model accuracy and adversely affects user experience. Conversely, for high-risk queries, an insufficient noise level may not sufficiently mask membership information, leaving the model vulnerable to membership inference attacks.
The central challenge, therefore, is to balance the trade-off between privacy and utility. A static noise injection strategy may either compromise model performance or inadequately protect against adversaries, thereby necessitating a more refined approach that dynamically adapts to the varying sensitivity across different queries.
@@ 4. Threat Model
In our threat model, the adversary has black-box access to the target model. That is, the attacker can submit any input query $q$ and obtain its corresponding prediction vector $f(q)$. The adversary is assumed to know the input/output format (for example, the number of classes and the range of confidence values) and may either have knowledge of the model’s architecture and training algorithm or only interact with the model via a machine-learning-as-a-service (MLaaS) platform where such internal details are hidden.
The adversary’s goal is to infer whether a given data record $q$ was used in training the target model. Formally, the attacker aims to distinguish between the hypotheses:
$H_{0}:\quad q\notin D_{\text{train}},$
$H_{1}:\quad q\in D_{\text{train}}.$
Where $D_{\text{train}}$ stands for the training dataset. We assume that the adversary may also have some background knowledge about the population from which $D_{\text{train}}$ is drawn, such as general statistics on feature distributions. An attack is considered successful if the adversary can determine the membership status of $q$ with high accuracy.
Our proposed defense mechanism, DynaNoise, is designed under these assumptions. By dynamically adjusting the noise added to the model’s outputs based on the sensitivity of each query, DynaNoise aims to blur the distinction between training and non-training data, thereby reducing the adversary’s ability to reliably infer membership.

### v2

# Dynamic Probabilistic Noise Injection for Membership Inference Defense

@@ Abstract.
Membership Inference Attacks (MIAs) expose privacy risks by determining whether a specific sample was part of a model’s training set. These threats are especially serious in sensitive domains such as healthcare and finance. Traditional mitigation techniques, such as static differential privacy, rely on injecting a fixed amount of noise during training or inference. However, this often leads to a detrimental trade-off: the noise may be insufficient to counter sophisticated attacks or, when increased, can substantially degrade model accuracy. To address this limitation, we propose DynaNoise, an adaptive inference-time defense that modulates injected noise based on per-query sensitivity. DynaNoise estimates risk using measures such as Shannon entropy and scales the noise variance accordingly, followed by a smoothing step that re-normalizes the perturbed outputs to preserve predictive utility. We further introduce MIDPUT (Membership Inference Defense Privacy-Utility Trade-off), a scalar metric that captures both privacy gains and accuracy retention. Our evaluation on several benchmark datasets demonstrates that DynaNoise substantially lowers attack success rates while maintaining competitive accuracy, achieving strong overall MIDPUT scores compared to state-of-the-art defenses.
membership inference attack, adaptive noise injection, privacy-preserving machine learning.
copyright: nonejournalyear: YYYYjournalvolume: YYYYjournalnumber: Xdoi: XXXXXXX.XXXXXXXconference: Proceedings on Privacy Enhancing Technologies; printacmref=false,printccs=false,printfolios=true
@@ 1. Introduction
Machine learning has revolutionized many domains by leveraging vast amounts of data to achieve impressive performance (Islam et al., 2024; Forough et al., 2023; Rayed et al., 2024; Zhang et al., 2019; Forough et al., 2022; Sahu et al., 2023). However, this success comes at a cost, where sensitive information from training datasets may be inadvertently memorized, posing serious privacy risks (Schwarzschild et al., 2024; El Mestari et al., 2024; Yang et al., 2024; Ye, 2024; Li et al., 2024). For example, in a scenario where a hospital deploys a predictive model to diagnose diseases; if an attacker can determine whether a patient’s record was part of the training set, it may reveal that the patient has visited the hospital, thereby compromising their privacy. Similarly, in finance, membership leakage could expose clients’ investment histories. Such risks underscore the need for robust defenses against privacy attacks.
Membership Inference Attacks (MIAs), as shown in Figure 1, exploit subtle differences in a model’s output behavior to determine whether a specific data record was used during training, thereby threatening the confidentiality of individual data points (Shokri et al., 2017). Conventional defenses, such as differential privacy, introduce a fixed level of noise during training or inference to obscure membership information (Abadi et al., 2016). While these approaches offer formal privacy guarantees, they force an inherent trade-off between privacy and utility. In many practical settings, uniformly adding noise can significantly degrade model performance, yet reducing the noise level may leave the model susceptible to advanced membership inference attacks.
Figure 1. Illustration of the membership inference attack (MIA) process.
To address these challenges, we introduce DynaNoise, an adaptive noise injection approach that modulates the privacy noise dynamically based on query sensitivity. Our method first assesses the risk of each query through sensitivity analysis, by utilizing metrics such as Shannon entropy (Sepúlveda-Fontaine and Amigó, 2024), and then adjusts the noise variance accordingly. A subsequent probabilistic smoothing step is applied to re-normalize the perturbed outputs, ensuring that the model retains high predictive accuracy while effectively obfuscating membership signals. Another key novelty of our approach is the introduction of a new empirical metric, the Membership Inference Defense Privacy-Utility Trade-off (MIDPUT), which quantitatively captures the balance between the reduction in attack success rates and the preservation of model accuracy. Our experimental results on CIFAR-10, ImageNet-10, and SST-2 datasets demonstrate that DynaNoise not only significantly reduces success rates of membership inference attacks but also achieves notably higher improvement in the MIDPUT metric compared to existing defenses.
@@
    Proposal of a novel post-hoc defense mechanism against membership inference attacks called DynaNoise that dynamically adjusts the noise level based on query sensitivity, thereby providing stronger privacy protection with minimal impact on target model accuracy.
@@
    Proposal of a new empirical metric called MIDPUT, that quantifies the trade-off between the reduction in attack success rates and the loss in model performance. This metric enables a more detailed and precise evaluation of privacy defenses.
@@
    Comprehensive experimental study on multiple benchmarks (CIFAR-10, ImageNet-10, and SST-2), systematically varying DynaNoise parameters and benchmarking against state-of-the-art defenses (AdvReg, MemGuard, RelaxLoss, SELENA, and HAMP). We also evaluate robustness under strong membership inference attacks, including LiRA, ensuring a fair and rigorous assessment.
@@ 2.1. Membership Inference Attacks (MIAs)
Shadow training-based attacks. Shokri et al. (Shokri et al., 2017) introduced shadow-model MIAs, where an adversary trains auxiliary models on data from a similar distribution and uses their outputs to train an attack classifier distinguishing members from non-members. Salem et al. (Salem et al., 2018) showed that this approach can be effective even with a single shadow model, substantially reducing the attacker’s cost. Long et al. (Long et al., 2020) further improved practicality by refining shadow-model construction and the attack classifier to reduce query requirements. Overall, shadow-model attacks remain strong but typically rely on access to representative auxiliary data from the target distribution.
Metric-based attacks. Metric-based attacks infer membership directly from the target model’s outputs by evaluating simple statistics, without training auxiliary models. Yeom et al. (Yeom et al., 2018) propose a loss-based attack that exploits the tendency of models to incur lower loss on training samples, inferring membership via thresholding. Song et al. (Song and Mittal, 2021) similarly introduce confidence-based attacks that predict membership based on whether the maximum confidence exceeds a predefined threshold. While lightweight and effective, these methods are sensitive to threshold selection and model calibration.
Song and Mittal (Song and Mittal, 2021) further extend this line of work by proposing entropy-based and modified entropy (M-Entropy) attacks, which exploit the observation that training samples typically produce lower-entropy and more structured output distributions. The entropy attack infers membership by thresholding Shannon entropy, while M-Entropy incorporates the probability assigned to the true label to improve robustness across confidence profiles. These attacks operate solely on output probabilities, making them well suited for black-box settings.
Likelihood Ratio Attack (LiRA). Carlini et al. (Carlini et al., 2022) recently introduced LiRA, which reframes membership inference as a statistical hypothesis test. Unlike earlier shadow-model or threshold-based approaches, LiRA estimates the distribution of model outputs on each example when it is included in training versus excluded. By fitting Gaussian distributions to these per-example losses using shadow models, LiRA computes a likelihood ratio to decide membership. This per-example calibration makes the attack far more effective at extremely low false-positive rates. Empirically, LiRA achieves up to 10x higher true-positive rates than prior attacks while also dominating in traditional aggregate metrics, establishing it as one of the strongest known MIAs.
@@ 2.2. Membership Inference Defenses
Abadi et al. (Abadi et al., 2016) propose DP-SGD, which enforces differential privacy during training by clipping per-example gradients and adding calibrated Gaussian noise, with a moments accountant used to track cumulative privacy loss. This provides formal, provable privacy guarantees at the model level. However, DP-SGD applies a static noise injection scheme that does not adapt to the sensitivity of individual queries, and the large noise required to satisfy strict privacy budgets often leads to significant accuracy degradation, particularly for complex or high-dimensional models.
Nasr et al. (Nasr et al., 2018) introduce adversarial regularization, a training-time defense that formulates membership inference mitigation as a min–max game between a classifier and an auxiliary attack model. By jointly training the classifier to minimize attack success, this approach can reduce membership leakage with modest utility loss, but it provides no formal privacy guarantees and is sensitive to hyperparameter tuning and dataset characteristics. In a related but complementary direction, Chen et al. (Chen et al., 2022) propose RelaxLoss, which mitigates membership inference by reducing the train–test loss gap through loss relaxation and posterior flattening during training. By smoothing overconfident predictions, RelaxLoss effectively lowers attack accuracy while largely preserving utility; however, it operates solely at training time and requires retraining the model from scratch, limiting its applicability as a post-hoc defense.
@@
Finally, Tang et al. (Tang et al., 2022) take an ensemble-based approach with SELENA, which trains multiple sub-models on overlapping subsets of the training data and then distills their outputs into a single prediction through a self-distillation process. This adaptive inference strategy selectively aggregates predictions from sub-models that have not seen the queried sample, thereby enhancing the privacy-utility trade-off. Despite its advantages, the ensemble inference process introduces moderate computational overhead due to the requirement of running multiple sub-models concurrently in the training phase, which can be a drawback in resource-constrained settings.
In a complementary line of work, Chen et al. (Chen and Pattabiraman, 2024) propose HAMP, which mitigates membership inference by reducing model overconfidence. HAMP achieves this through a training framework that employs high-entropy soft labels and an entropy-based regularizer to discourage overly confident predictions. In addition, HAMP applies a testing-time output modification that uniformly converts prediction scores into low-confidence distributions while preserving the predicted label. As a result, confidence suppression is applied in a query-independent manner at inference time. While this approach can effectively reduce attack success rates, its uniform suppression may unnecessarily perturb low-risk predictions, leading to noticeable degradation in target model accuracy, particularly under strong membership inference attacks.
@@
Novelty compared to prior noise-based defenses: Prior defenses against membership inference attacks have perturbed model outputs with random noise, either during training or at inference time, typically using constant magnitudes or coarse, confidence-based heuristics, such as relying solely on the highest predicted probability, to set the noise level. While such approaches can reduce attack success rates, they often fail to fully leverage the structure of the entire output distribution and may apply suboptimal noise amounts across queries. In addition, defenses such as HAMP (Chen and Pattabiraman, 2024) mitigate membership leakage by reducing overconfidence and applying a uniform, query-independent confidence suppression at inference time, without adapting the level of protection to the uncertainty or risk of individual queries.
DynaNoise departs from these patterns in several key ways. First, rather than basing noise magnitude solely on the highest predicted probability or applying uniform suppression, we derive a query-specific sensitivity score from the Shannon entropy of the full logit distribution, capturing the model’s overall uncertainty. Second, we integrate Gaussian perturbation with a principled temperature-scaling step to re-normalize the perturbed logits, ensuring controlled smoothing of the outputs. Third, our adaptation is continuous and fine-grained, with the noise level varying smoothly across queries instead of being assigned via thresholds or bins. Finally, DynaNoise is applied purely as a post-hoc mechanism, requiring no model retraining or architectural modifications, making it lightweight and easily deployable in practice.
@@
Defense Method Year Approach Privacy Guarantee Utility Impact Comp. Overhead
@@
MemGuard (Jia et al., 2019) 2019 Post-processing: adversarial noise added to confidence score vectors Empirical Minimal accuracy drop High (Inference)
RelaxLoss (Chen et al., 2022) 2022 Loss-relaxation with posterior flattening and alternating ascent/descent to reduce train–test loss gap (confidence smoothing) Empirical Low to moderate accuracy drop Low
@@
HAMP (Chen and Pattabiraman, 2024) 2024 Training-time overconfidence mitigation with entropy regularization and high-entropy soft labels, combined with uniform confidence suppression at inference time Empirical Moderate to high accuracy drop Low
DynaNoise (This work) 2026 Adaptive noise injection based on query sensitivity (sensitivity analysis, dynamic noise variance modulation, probabilistic smoothing) Empirical Minimal accuracy drop Low
@@ 3. Problem Statement
Let $f_{\theta}:\mathcal{X}\rightarrow\mathcal{Y}$ denote a trained model parameterized by $\theta$, which outputs a probability distribution $f_{\theta}(q)=(p_{1},p_{2},\ldots,p_{k})$ over $k$ classes for an input sample $q\in\mathcal{X}$. Membership inference attacks aim to determine whether a given query $q$ was part of the model’s training dataset $D_{\text{train}}$. Formally, an adversary attempts to infer the membership status of $q$ by distinguishing between two hypotheses:
$H_{0}:q\notin D_{\text{train}},\qquad H_{1}:q\in D_{\text{train}}.$
Such attacks exploit the distributional discrepancies between the model’s output probabilities for members versus non-members, i.e., $f_{\theta}(q_{m})\neq f_{\theta}(q_{n})$ for $q_{m}\in D_{\text{train}}$ and $q_{n}\notin D_{\text{train}}$.
Traditional privacy-preserving mechanisms, such as differential privacy (DP), mitigate this risk by adding a fixed amount of random noise $\eta$ to the model outputs:
$\tilde{f}_{\theta}(q)=f_{\theta}(q)+\eta,\quad\eta\sim\mathcal{N}(0,\sigma^{2}I),$
where $\sigma$ is a pre-defined noise scale. However, this uniform perturbation fails to consider the heterogeneous sensitivity of queries. For low-risk queries, excessive noise can unnecessarily degrade prediction accuracy, while for high-risk queries, insufficient noise may still leak membership information.
The key challenge, therefore, is to design a mechanism that achieves an optimal balance between privacy and utility by adapting the injected noise to the sensitivity of each query. Specifically, the goal is to learn or define a sensitivity function $R(q)$ such that the noise variance $\sigma^{2}(q)$ scales with the privacy risk of the query, thereby maximizing protection for sensitive cases without impairing overall model performance.
@@ 4. Threat Model
In our threat model, the adversary is granted black-box access to the target model. Specifically, the attacker can submit any input query $q$ and obtain its corresponding prediction vector, i.e., the post-softmax probability output $\hat{f}(q)$. The adversary is assumed to know the input/output format (e.g., number of classes and range of confidence values) and may either (i) possess partial knowledge of the model architecture and training procedure, or (ii) interact with the model through a machine-learning-as-a-service (MLaaS) interface where internal parameters remain hidden. We assume that the adversary has access to the full prediction vector (e.g., confidence or probability scores). As a result, DynaNoise is designed to defend against score-based membership inference attacks and does not address attacks that rely solely on predicted labels.
The adversary’s objective is to determine whether a given record $q$ was used in training the target model. Formally, the attack attempts to distinguish between:
$H_{0}:\;q\notin D_{\text{train}},\qquad H_{1}:\;q\in D_{\text{train}},$
where $D_{\text{train}}$ denotes the model’s training dataset. The adversary may also have access to auxiliary information about the underlying data distribution, such as population-level feature statistics. An attack is considered successful if it can infer the membership status of $q$ with accuracy significantly better than random guessing.
Query assumptions: DynaNoise operates under a limited-query assumption, where repeated identical queries are not allowed or are infrequent. This reflects common practical scenarios in deployed MLaaS systems that log or rate-limit queries to prevent excessive access. If the same query were submitted multiple times, an adversary could theoretically average the noisy outputs and approximate the unperturbed logits. Addressing such adaptive query repetition is outside the current scope but can be mitigated by caching consistent noisy responses or incorporating query-count–dependent noise scaling, as discussed in Section 7.6.
Under these conditions, DynaNoise dynamically adjusts the noise added to each output based on the sensitivity of the query, thereby reducing the adversary’s ability to distinguish between members and non-members while preserving model utility.